# On the existence of unbiased resilient estimators in discrete quantum systems

### v3

@@ V Conclusions
Since long ago, it has been clear in estimation theory that while the Cramér-Rao bound (CRB) is extremely useful and illustrative, it is not always meaningful. For that reason, different bounds were constructed and introduced in the classical regime. On the other hand, until very recently, there were no works extending these bounds to the quantum regime [30].
@@
Moreover, following a similar reasoning, we could anticipate when computing higher orders in the BhB gives no extra information. Specifically, we saw that for $N$ outcomes, it is “sufficient” to calculate the $n$-th order BhB bound where $n\leq N-1$. However, in the quantum case, where we have more freedom in the measurements, the order is lifted to $n\leq\frac{N(N+1)}{2}-1$, implying that in the quantum regime, the estimators are more resilient. Finally, we compute the Bhattacharyya bound for a Mach-Zehnder interferometer and numerically show that this bound scales inversely proportional to the square of the number of photons, which is the Heisenberg scaling. Since Bhattacharyya bounds are tighter, this result supports the claim that it is possible to achieve Heisenberg scaling in the Mach-Zehnder interferometer using a more robust estimator.
We note that the existence of quantum Barankin bounds for finite dimensional systems has also been studied recently in Ref. [39].
@@ Appendix A Existence of the estimator
We have proved that BhB is unbounded if and only if Eq. (38) is true. We note that Eq. (38) is the same as Eq. (30). This is the same as saying that there is no estimator satisfying the conditions. As stated before, the matrix of Eq. (24) has at most $N$ linearly independent rows associated with $N$ independent conditions. Consider the case where the first $N-1$ higher order derivatives are linearly independent, then $n^{th}$order BhB bound will be equal to the $(N-1)^{th}$ BhB bound if $n\geq N-1$.
Appendix B Example of the quantum case
In this section, we show a case in which we see the lifting in the constraints produced by the quantum nature of the problem, meaning that we see that the quantum bounds are not trivial –giving the same value as bounds from lower orders– for $n>N-1$, being $n$ the order of the BhB employed and $N$ the number of measurement outcomes. As in this example we deal with qubits, $N=2$.
$\rho(0)=\frac{1}{2}\mathbb{I}+\frac{(2\lambda-1)}{2}\sigma_{z},$ (39)
where $\lambda\in(0,1)$ and
where the first element corresponds to the QFI. To finally get the BhB, we take the inverse of the matrix and look at the first element of it. In Fig. (5) we can see the lower bound on the variance of the estimator when using the QCR bound and QBhB. Thus, computing the quantum Bhattacharyya bound gives new non-trivial information even though $n>N-1$.
Figure 5: QCR and QBhB as a function of $\lambda$. One sees that they differ and as expected QBhB bound will be higher than QCR. In addition, we plot these values of $\lambda$ because the plot is symmetric around $\lambda=0.5$ and also divergent for this same value –since the state will be $\rho\propto\mathbb{I}$ and won’t contain any information on the parameter–.
$\alpha(Q^{-1}\bm{\lambda})\bm{L},$
$P_{\theta}^{\text{CRB}}(\alpha_{2})$ $=1-P_{\theta}^{\text{CRB}}(\alpha_{1})$
where $f=2(\theta-\phi)(\theta+\phi)$ and $\alpha_{i}$ are the eigenvalues of $L_{1}$.
The probability distribution that saturates the BhB is,
$P_{\theta}^{Bh}(\beta_{2})=$ $1-P_{\theta}^{Bh}(\beta_{1})$
and

### v4

@@ V Conclusions
Since long ago, it has been clear in estimation theory that while the Cramér-Rao bound (CRB) is extremely useful and illustrative, it is not always meaningful. For that reason, different bounds were constructed and introduced in the classical regime. On the other hand, until very recently, there were no works extending these bounds to the quantum regime [30, 31].
@@
Moreover, following a similar reasoning, we could anticipate when computing higher orders in the BhB gives no extra information. Specifically, we saw that for $N$ outcomes, it is “sufficient” to calculate the $n$-th order BhB bound where $n\leq N-1$. However, in the quantum case, where we have more freedom in the measurements, the order is lifted to $n\leq\frac{N(N+1)}{2}-1$, implying that in the quantum regime, one can construct tighter bounds, i.e. higher order BhB.
Finally, we compute the Bhattacharyya bound for a Mach-Zehnder interferometer and numerically show that this bound scales inversely proportional to the square of the number of photons, which is the Heisenberg scaling. Since Bhattacharyya bounds are tighter, this result supports the claim that it is possible to achieve Heisenberg scaling in the Mach-Zehnder interferometer using a more robust estimator.
@@ Appendix A Existence of the estimator
We have proved that BhB is unbounded if and only if Eq. (44) is true. We note that Eq. (44) is the same as Eq. (36). This is the same as saying that there is no estimator satisfying the conditions. As stated before, the matrix of Eq. (III.1) has at most $N$ linearly independent rows associated with $N$ independent conditions. Consider the case where the first $N-1$ higher order derivatives are linearly independent, then $n^{th}$order BhB bound will be equal to the $(N-1)^{th}$ BhB bound if $n\geq N-1$.